OpenStudy (anonymous):

A principal of $10,000 is invested in an account paying an annual percentage rate of 5%. Round all dollar amounts to two decimal places. Find the amount in the account after 10 years if the account is compounded n times a year. n= 4

OpenStudy (wolf1728):

The formula is located here: http://www.1728.org/compint3.htm Total = principal * [1 + (rate/n)]^n * years where 'n' is the number of compounding periods per year Total = 10,000 * [1 + (.05/4)]^4 * 10 Total = 10,000 * 1.6436194635 Total = 16,436.19 Here's a calculator to check your work: http://www.1728.org/compint.htm